# Quench Spectroscopy for Dissipative and Non-Hermitian Quantum Lattice Models

### v1

@@ 7 Applicability of the quench spectroscopy to a non-Hermitian quantum lattice model
where $\hat{S}^{\alpha}_{R}$ corresponds to the $s=1/2$ spin operator acting on the lattice site $R$ along the $\alpha\in\{x,z\}$ axis. $J>0$ refers to the antiferromagnetic spin exchange coupling between nearest neighbors spins, $h>0$ is the constant and homogeneous transverse magnetic field and $\gamma$ denotes the dissipative strength representing the local dephasing noise. In what follows, we consider $\gamma>0$ as well as the paramagnetic phase (also called $z$ polarized phase in the literature) where the spins are polarized along $z$ axis, i.e. are aligned with the magnetic field, for $h\gg J$. In Ref. [74], J.Despres et al. studied the quasiparticle dispersion relation as well as the quench dynamics of this non-Hermitian spin model in the paramagnetic phase. Regarding the low-lying excitation spectrum, they relied on the following Holstein-Primakoff (HP) transformation:
$\hat{S}^{x}_{R}=\frac{\hat{a}_{R}+\hat{a}^{{\dagger}}_{R}}{2};$ (46a)
$\hat{S}^{y}_{R}=\frac{\hat{a}_{R}-\hat{a}^{{\dagger}}_{R}}{2i};$ (46b)
$\hat{S}^{z}_{R}=\frac{1}{2}-\hat{a}^{{\dagger}}_{R}\hat{a}_{R},$ (46c)
@@
To investigate the quench dynamics of the non-Hermitian TFIM, J.Despres et al. derived the EoM associated to the quadratic bosonic correlator $G_{k}(t)=\langle\hat{a}^{{\dagger}}_{k}\hat{a}_{k}\rangle_{t}$ and $F_{k}(t)=\langle\hat{a}_{k}\hat{a}_{-k}\rangle_{t}$ where $\langle...\rangle_{t}=\langle\Psi(t)|...|\Psi(t)\rangle$. The time-evolved many-body quantum state |$\ket{\Psi(t)}$ reads:
|$\ket{\Psi(t)}=\frac{e^{-i\hat{H}t}\ket{\Psi_{0}}}{||e^{-i\hat{H}t}\ket{\Psi_{0}}||},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \ket{\Psi_{0}}=\ket{\Psi(0)}=\ket{\mathrm{GS}(\hat{H}(\gamma=0))},$ (47)
where |$\ket{\Psi_{0}}$ denotes the initial many-body quantum state corresponding to the ground state of the Hermitian version of the Hamiltonian $\hat{H}$, i.e. $\gamma=0$. Using the latter expression, the EoM associated to the time-dependent expectation value of any observable $\hat{A}$ denoted by $G_{A}(t)=\langle\Psi(t)|\hat{A}|\Psi(t)\rangle=\langle\hat{A}\rangle_{t}$ reads as:
@@
$\frac{\mathrm{d}}{\mathrm{d}t}F_{k}(t)=4\operatorname{Im}(\mathcal{A}_{k})F_{k}(t)G_{k}(t)-2i\mathcal{A}_{k}F_{k}(t)-i\mathcal{B}_{k}\left[1+2G_{k}(t)\right];$ (49a)
$\frac{\mathrm{d}}{\mathrm{d}t}G_{k}(t)=-2\mathcal{B}_{k}\operatorname{Im}(F_{k}(t))+2\operatorname{Im}(\mathcal{A}_{k})\left(|F_{k}(t)|^{2}+G_{k}(t)+G_{k}(t)^{2}\right),$ (49b)
Figure 7: (Left) $G_{1}(R,t)=\langle\hat{a}^{{\dagger}}_{R}\hat{a}_{0}\rangle_{t}$ the equal-time one-body correlation function (right) the associated quench spectral function $S_{k}(\omega)$ for a weak sudden global quench on $\gamma$ from $\gamma=0$ to $\gamma>0$ for the non-Hermitian 1D TFIM confined in the paramagnetic ($z$ polarized) phase. The dashed-dotted magenta lines represent twice and minus twice the real part of the momentum-dependent function $\mathcal{A}_{k}$, i.e. $2\operatorname{Re}(\mathcal{A}_{k})$ and $-2\operatorname{Re}(\mathcal{A}_{k})$ respectively. For clarity, the space-time correlation pattern of $G_{1}$ for the calculation of $S_{k}(\omega)$ is plotted for $R\geq 0$ and $R\leq 0$ and $t\geq 0$ to recover the second half of the effective quasiparticle dispersion relation. The parameters are: $L=400$ to reach the thermodynamic limit, $J=1$, $h=3.5$ and $\gamma=0.02$.
@@
$F_{k}(0)=\cosh(\alpha_{k})\sinh(\alpha_{k});$ (50)
$G_{k}(0)=\sinh^{2}(\alpha_{k}),$ (51)
$\alpha_{k}=\frac{1}{2}\mathrm{arctanh}\left(-\frac{\mathcal{B}_{k}}{\mathcal{A}_{k}(\gamma=0)}\right).$ (52)
Using the set of EoMs defined at Eq. (49), the correlation spreading induced by the quench on $\gamma$ is investigated in Ref. [74]by analyzing the equal-time one-body correlation function $G_{1}(R,t)$ which reads as:
@@
$G_{1}(R,t)=\frac{1}{L}\sum_{k}\mathcal{F}_{k}^{(1)}\cos(kR)\cos(2\operatorname{Re}(\mathcal{A}_{k})t)\exp(2\operatorname{Im}(\mathcal{A}_{k})t),$ (54)
@@
$G_{1}(R,t)\sim\int_{\mathcal{B}}\mathrm{d}k\leavevmode\nobreak\ \mathcal{S}_{k}^{(1)}\left\{e^{i(kR+2\operatorname{Re}(\mathcal{A}_{k})t)}+e^{i(kR-2\operatorname{Re}(\mathcal{A}_{k})t)}\right\}.$ (55)

### v2

@@ IV Applicability of the quench spectroscopy to a non-Hermitian quantum lattice model
where $\hat{S}^{\alpha}_{R}$ corresponds to the $s=1/2$ spin operator acting on the lattice site $R$ along the $\alpha\in\{x,z\}$ axis. $J>0$ refers to the antiferromagnetic spin exchange coupling between nearest neighbors spins, $h>0$ is the constant and homogeneous transverse magnetic field and $\gamma$ denotes the dissipative strength representing the local dephasing noise. In what follows, we consider $\gamma>0$ as well as the paramagnetic phase (also called $z$ polarized phase in the literature) where the spins are polarized along $z$ axis, i.e. are aligned with the magnetic field, for $h\gg J$. In Ref. despres2024 , the Authors studied the quasiparticle dispersion relation as well as the quench dynamics of this non-Hermitian spin model in the paramagnetic phase. Regarding the low-lying excitation spectrum, they relied on the following Holstein-Primakoff (HP) transformation:
$\hat{S}^{x}_{R}=\frac{\hat{a}_{R}+\hat{a}^{{\dagger}}_{R}}{2},\leavevmode\nobreak\ \hat{S}^{y}_{R}=\frac{\hat{a}_{R}-\hat{a}^{{\dagger}}_{R}}{2i},\leavevmode\nobreak\ \hat{S}^{z}_{R}=\frac{1}{2}-\hat{a}^{{\dagger}}_{R}\hat{a}_{R},$ (39)
@@
To investigate the quench dynamics of the non-Hermitian TFIM, the Authors of Ref. despres2024 derived the EoM associated to the quadratic bosonic correlator $G_{k}(t)=\langle\hat{a}^{{\dagger}}_{k}\hat{a}_{k}\rangle_{t}$ and $F_{k}(t)=\langle\hat{a}_{k}\hat{a}_{-k}\rangle_{t}$ where $\langle...\rangle_{t}=\langle\Psi(t)|...|\Psi(t)\rangle$. The time-evolved many-body quantum state |$\ket{\Psi(t)}$ is defined as:
|$\ket{\Psi(t)}=\frac{e^{-i\hat{H}t}\ket{\Psi_{0}}}{||e^{-i\hat{H}t}\ket{\Psi_{0}}||},$ (40)
where |$\ket{\Psi_{0}}=\ket{\Psi(0)}=\ket{\mathrm{GS}(\hat{H}(\gamma=0))}$ denotes the initial many-body quantum state corresponding to the ground state of the Hermitian version of the Hamiltonian $\hat{H}$ implying $\gamma=0$. Using the latter expression, the EoM associated to the time-dependent expectation value of any observable $\hat{A}$ denoted by $G_{A}(t)=\langle\Psi(t)|\hat{A}|\Psi(t)\rangle=\langle\hat{A}\rangle_{t}$ reads as:
@@
$\frac{\mathrm{d}}{\mathrm{d}t}G_{k}(t)=$ $\leavevmode\nobreak\ -2\mathcal{B}_{k}\operatorname{Im}(F_{k}(t))+2\operatorname{Im}(\mathcal{A}_{k})\left(|F_{k}(t)|^{2}+G_{k}(t)+G_{k}(t)^{2}\right);$ (42a)
$\frac{\mathrm{d}}{\mathrm{d}t}F_{k}(t)=$ $\leavevmode\nobreak\ 4\operatorname{Im}(\mathcal{A}_{k})F_{k}(t)G_{k}(t)-2i\mathcal{A}_{k}F_{k}(t)-i\mathcal{B}_{k}\left[1+2G_{k}(t)\right],$ (42b)
Figure 7: (Top panel) $G_{1}(R,t)=\langle\hat{a}^{{\dagger}}_{R}\hat{a}_{0}\rangle_{t}$ the equal-time one-body correlation function (bottom panel) the associated quench spectral function $S_{k}(\omega)$ for a weak sudden global quench on $\gamma$ from $\gamma=0$ to $\gamma>0$ for the non-Hermitian 1D TFIM confined in the paramagnetic ($z$ polarized) phase. The dashed-dotted magenta lines represent twice and minus twice the real part of the momentum-dependent function $\mathcal{A}_{k}$, i.e. $2\operatorname{Re}(\mathcal{A}_{k})$ and $-2\operatorname{Re}(\mathcal{A}_{k})$ respectively. For clarity, the space-time correlation pattern of $G_{1}$ for the calculation of $S_{k}(\omega)$ is plotted for $R\geq 0$ and $R\leq 0$ and $t\geq 0$ to recover the second half of the effective quasiparticle dispersion relation. The dimensionless transverse-field and dissipation parameters are given by $h/J=3.5$ and $\gamma/J=0.02$ respectively.
@@
$F_{k}(0)=\cosh(\alpha_{k})\sinh(\alpha_{k});$ (43a)
$G_{k}(0)=\sinh^{2}(\alpha_{k}),$ (43b)
$\alpha_{k}=\frac{1}{2}\mathrm{arctanh}\left(-\frac{\mathcal{B}_{k}}{\mathcal{A}_{k}(\gamma=0)}\right).$ (43c)
Using the set of EoMs defined at Eq. (42), the correlation spreading induced by the quench on $\gamma$ is investigated in Ref. despres2024 by analyzing the equal-time one-body correlation function $G_{1}(R,t)$ which reads as:
@@
$G_{1}(R,t)=\frac{1}{L}\sum_{k}\mathcal{F}_{k}^{(1)}\cos(kR)\cos[2\operatorname{Re}(\mathcal{A}_{k})t]e^{2\operatorname{Im}(\mathcal{A}_{k})t},$ (45)
@@
$G_{1}(R,t)\sim\int_{\mathcal{B}}\mathrm{d}k\mathcal{S}_{k}^{(1)}\left\{e^{i[kR+2\operatorname{Re}(\mathcal{A}_{k})t]}+e^{i[kR-2\operatorname{Re}(\mathcal{A}_{k})t]}\right\}.$ (46)